### 22 Dec 2023 v1

# Pub/Sub Message Brokers for GenAI

Abstract
In today’s digital world, Generative Artificial Intelligence (GenAI) such as Large Language Models (LLMs) is becoming increasingly prevalent, extending its reach across diverse applications. This surge in adoption has sparked a significant increase in demand for data-centric GenAI models, highlighting the necessity for robust data communication infrastructures. Central to this need are message brokers, which serve as essential channels for data transfer within various system components. This survey aims to delve into a comprehensive analysis of traditional and modern message brokers, offering a comparative study of prevalent platforms. Our study considers numerous criteria including, but not limited to, open-source availability, integrated monitoring tools, message prioritization mechanisms, capabilities for parallel processing, reliability, distribution and clustering functionalities, authentication processes, data persistence strategies, fault tolerance, and scalability. Furthermore, we explore the intrinsic constraints that the design and operation of each message broker might impose, recognizing that these limitations are crucial in understanding their real-world applicability. We then leverage these insights to propose a sophisticated message broker framework—one designed with the adaptability and robustness necessary to meet the evolving requisites of GenAI applications. Finally, this study examines the enhancement of message broker mechanisms specifically for GenAI contexts, emphasizing the criticality of developing a versatile message broker framework. Such a framework would be poised for quick adaptation, catering to the dynamic and growing demands of GenAI in the foreseeable future. Through this dual-pronged approach, we intend to contribute a foundational compendium that can guide future innovations and infrastructural advancements in the realm of GenAI data communication.
keywords:
Generative AI, Message Brokers, Publish/Subscribe Paradigm, Brokerless, MLOps, Continuous Diagnostics and Mitigation, Large Language Models.
1 Introduction
In the burgeoning field of Generative Artificial Intelligence (GenAI), the computing continuum faces unprecedented challenges in efficiently managing data flows and computational resources. GenAI has predominantly targeted consumer applications, offering applications such as the chatGPT, based on a large language model (LLMs), a type of machine learning (ML) model with a deep neural network. However, a surge in machine-to-machine (M2M) use cases beckons a reevaluation of the supporting communication infrastructure. This survey paper delves into the role of publish/subscribe (pub/sub) message broker systems, considering in particular their emerging role for seamless and scalable data exchange in GenAI applications. We scrutinize contemporary message brokers for their adaptability and efficiency in GenAI contexts, outline existing challenges, and chart promising research avenues for future development.
In more detail, while the focus of GenAI systems has been on consumer-oriented applications, there is an increased interest towards M2M use cases. GenAI has been proposed to be used in, for example, for networking, wireless communication, and compression bariah2023large .
As a result, current computing continuum platforms, spanning the networks and computational resources from user devices to cloud kokkonen2022autonomy ; motlagh2022edge , face new challenges. These platforms provide support AI models, offering interconnect between their data sources and sinks, and optimising their use of resources in the computing continuum. As GenAI models require and generate ever larger amounts of data wang2023overview , all the while consuming computational resources varying from moderate to massive kokkonen2022autonomy , the computing continuum must offer a dynamic and scalable communication and computation substrate to ensure timely data dissemination and efficient use of resources dustdar2022distributed ; computers12100198 .
Approaches based on the publish/subscribe paradigm (in short, pub/sub) are a preferred choice for a variety of real-world AI applications in the computing continuum. Pub/sub approaches decouple data producers from their consumers, allowing applications to develop components independently, and enhancing system robustness and adaptability eugster2003many . Furthermore, pub/sub makes system design more flexible by increasing the independence between system components with a reliable interconnect.
Pub/sub systems are based on the exchange of data between clients (e.g., services or application components) through a message broker. Publishers submit content to the broker, which then allows subscribers to access that content without knowing its source tarkoma2012publish ; donta2022survey . By managing, filtering, and routing communication between publishers and subscribers, the message broker acts as an intermediary layer hasenburg2020managing , routing and distributing messages efficiently, accurately, and in a timely manner, based on the interests expressed by subscribers. Within the brokers, message queues can temporarily store the messages, protecting the system from overflows or outages. Moreover, brokers also often provide other essential features such as persistent storage, monitoring, and authentication.
With the advent and growing popularity of GenAI, there is an increased need for robust and adaptive message brokers, equipped to cater to the specific requirements of GenAI applications. Such brokers must be capable of handling the enormous amounts of data required and generated by GenAI models, as well as adeptly monitoring the processes and the data exchanges for high Quality of Service (QoS) loven2023can .
The primary contributions of this survey are summarized as follows:
  • We provide a comprehensive review of recent message brokers, evaluating the brokers by their suitability for GenAI systems. We aim to guide the development of a compatible brokering framework in consideration of the evolving requirements for GenAI systems in the future.
  • We summarize the challenges of message brokers and highlight the need for a robust and efficient data communications substrate based on an increase in GenAI applications.
  • We discuss central research topics and their potential focus areas for making message brokers suitable for GenAI applications. We also describe promising algorithms for implementing such brokers.
This paper’s remaining sections are organized as follows. Section 2 provides a general definition of pub/sub paradigm and highlights the advantages and disadvantages of both broker-based and brokerless messaging architectures. Section 3 presents existing message brokers with their features and cons. Section 4 examines possible ways to make message brokers suitable for GenAI applications. The paper concludes with Section 5.
2 The Pub/Sub Paradigm
Pub/sub is a messaging paradigm where publishers send messages without indicating specific recipients. Remaining oblivious to the original publishers, subscribers receive relevant messages according to their interests. At its core, pub/sub thus decouples message delivery from the senders and recipients. This enhances the system’s adaptability and robustness, as it allows publishers and subscribers to operate independently eugster2003many . Furthermore, subscribers can flexibly choose topics based on their interests, enabling them to find content relevant to their preferences. As a result of pub/sub, real-time messaging can be sent to a wide range of subscribers, enabling scalable and timely dissemination of information donta2022survey ; chafi2022introduction ; maniezzo2023self .
As part of the pub/sub communication model, a message broker functions as an intermediary layer, managing the flow of messages from publishers to subscribers. By ensuring that messages are accurately routed to subscribers, based on their expressed interests or specific topics, the broker ensures that messages are received by subscribers precisely as they have been requested. By providing a layer of abstraction between publishers and subscribers, the broker goes beyond simply facilitating message transmission. Therefore, neither party needs to be aware of the other’s operations or presence. A key strength of the broker is its reliability pedrosa2021reducing , as it has mechanisms for guaranteeing delivery of messages even when subscribers are temporarily offline or have connectivity difficulties.
Moreover, message queues are general-purpose components of the broker that temporarily store messages from publishers until they can be delivered to subscribers as part of the broker process john2017survey . By orchestrating the overall flow of messages, the message broker ensures that the appropriate subscribers receive the messages based on their subscriptions, while a message queue ensures that these messages are held and dispatched in an orderly manner, ensuring that publishers and subscribers are able to communicate in an asynchronously and decoupled ways as shown in Fig. 1.
Figure 1: The Publish/Subscribe paradigm.
2.1 Message broker development
Over the past 30 years, message broker technology has evolved and innovated significantly as shown in Fig. 2. With the development of message-oriented middleware (MOM) and the Java Message Service (JMS), the technology began to gain prominence between 1980 and 1999, as organizations began to require greater integration and communication linthicum2000enterprise ; perry2001mqseries . During the period 2000-2009, the technology underwent significant advancements, influenced by the implementation of service-oriented architecture (SOA), the increased internet usage, and the emergence of cloud computing hohpe2004enterprise . Among these developments was the implementation of web services standards and open-source alternatives, laying the foundation for future advances in cloud-based message brokers. During the following decade, from 2010 to 2019, the technology adapted to new demands such as real-time data processing, cloud computing, IoT, and microservices architectures john2017survey . Increasing demands for Internet of Things and real-time data processing have led to the rise of containerization platforms such as Docker and Kubernetes.
With the advent of cloud-native architectures, microservices, edge computing, and IoT demands, message broker technology continued to evolve unabated between 2020 and 2023 indrasiri2021design . The integration of AI and ML was particularly important for optimizing message routing, anomaly detection, and auto-scaling, addressing the complexities of growing data volumes. In the future, message broker technology will continue to evolve, leveraging advances in computing and communication. As 5G/6G technologies advance, cross-platform interoperability, and decentralized architectures are developed, a number of trends are set to shape its future. These include edge computing, quantum computing, enhanced security, serverless architectures, and advancements in the 5G/6G technologies.
Figure 2: The timeline of message broker evolution from 1990 to present.
2.2 Broker vs. Brokerless Messaging Architecture
A broker is the main unit for managing and monitoring data of pub/sub systems tarkoma2012publish , offering scalability, balanced load distribution, and optimal resource utilization, among others. Additionally, the message broker ensures that messages are reliably transmitted, preventing any loss of data.
Despite their benefits, certain challenges accompany message brokers, particularly concerning scalability and efficiency. As data volume grows, brokers face increased complexity, undermining scalability – a common issue also observed in brokerless pub/sub systems known for their simplicity, quick access, and improved efficiency. In such systems, publishers and subscribers interact directly, making discovery, management, and availability crucial factors. However, the absence of a central unit for overseeing message flows complicates supervision and control. Furthermore, this setup does not inherently guarantee reliable message delivery iakushkin2014messaging .
3 Survey of Message Brokers
In the past decade, numerous message brokers have been developed both in proprietary and open-source sectors. Each broker possesses unique features and pitfalls, influenced by their respective vendors and intended applications. This section focuses on the most commonly used message brokers, with their features and challenges concisely summarized in Tables 1, 2, 3 and 4. In following subsections, we categorize these message brokers based on their open-source availability and the priority-based delivery of messages (built-in priority-support) that ensures messages are delivered in priority order, with high-priority messages processed first.
3.1 Open Source Message Brokers
We found 30 message brokers that were available as open source. Out of these, 17 supported priority messages, while 13 did not. Each is discussed in more detail in below subsections.
3.1.1 Priority Support
Apache ActiveMQ ApacheActiveMQ is a Java-based message broker licensed under the Apache 2.0 license. Through the use of double layers of SSL/TLS security layers, it provides dual security levels. A distribution of Apache ActiveMQ provided by FuseSource, Fuse Message Broker Fuse supports J2EE integration capabilities such as Java Database Connectivity (JDBC), J2EE Connector Architecture (JCA), and Enterprise JavaBeans (EJB). Apache Qpid ApacheQpid , on the other hand, provides cloud-based messaging capabilities and supports queuing for structured message exchange, making it essential for distributed applications.
RabbitMQ RabbitMQ was developed by Rabbit Technologies Ltd in 2006 using the Erlang programming language and released under the Mozilla Public License. It supports multiple protocols, including AMQP, STOMP, and MQTT. HornetQ HornetQ , a Java application based on JBoss, provides a distributed messaging platform for enterprise-level applications using STOMP and AMQP protocols.
Red Hat AMQ RedHatAMQ is a messaging protocol based on Java for large-scale Internet business applications with no administrative costs, installation, or configuration required. Celery Celery is written in Python and supports multiple message brokers, including RabbitMQ and Redis. Using Java Messaging Service (JMS) API, JBoss Messaging JBossMessaging is a messaging broker provided by JBoss, a division of Red Hat for facilitating communication between different components or applications in a distributed system.
OpenMQ OpenMQ is implemented in Java and was developed by Oracle as an open source protocol. Beanstalk Beanstalk creates queues automatically with pure Python. Gearman Gearman is an optimized server written in C/C++ with a simple interface that provides low application overhead. Enduro/X Enduro/X is written in C and offers native APIs for C/C++. For enhanced interprocess communication, it utilizes in-memory POSIX kernel queues. As part of the WSO2 Integration platform, WSO2 Message Broker WSO2 is a message-based communication component.
HiveMQ HiveMQ is compatible with MQTTv3.1 and all subsequent versions. The Eclipse Public License (EPL) and Eclipse Distribution License (EDL) cover this implementation. Redis Redis is BSD-licensed, used by companies such as Uber, Instagram, and AirBNB for caching and messaging queues. With 100 million concurrent connections per cluster and sub-millisecond latency, EMQX EMQX can efficiently and reliably connect massive amounts of IoT devices. EMQX nodes can be bridged by other MQTT servers and cloud services to send messages across platforms. Additionally, it deploys and operates on all public cloud platforms. Apache Pulsar Pulsar is an open-source distributed messaging system developed as a queuing system, but it recently added event streaming features. It combines many Kafka and RabbitMQ features.
3.1.2 No Priority Support
Apache Kafka Kafka was developed by LinkedIn as a distributed streaming platform, supporting multiple data formats, including JSON, Avro, and XML. Furthermore, Java, Python, and Go are the official client libraries and several cloud platforms are supported, including Amazon Web Services, Microsoft Azure, and Google Cloud Platform. It provides a variety of tools for managing and monitoring Kafka clusters, such as Kafka Manager, Kafka Monitor, and Kafka Connect.
Apache RocketMQ ApacheRocketMQ is a cloud-native platform that operates across distributed systems, facilitating real-time data processing. With support for versions 5.0, 3.1.1, and 3.1, Eclipse Mosquitto EclipseMosquitto implements the MQTT protocol. ZeroMQ zeromq is supported by a large and active open source community, and utilizes a broker-less pub/sub pattern.
Apache NiFi ApacheNiFi , developed by the Apache Software Foundation, automates data exchange between software systems, and facilitates the conversion of data formats in real time. Ably Realtime AblyRealtime is built on Ably’s Data Stream Network, which includes a cloud network and realtime messaging fabric. Additionally, over 40 Client Library SDKs are available, as well as native support for six real-time protocols. Apache SamZa SamZa is a streaming framework based on Apache Kafka and Apache Hadoop developed by LinkedIn and now part of the Apache Software Foundation. It processes real-time data streams generated by Apache Kafka, Amazon Kinesis, and Azure Event Hub.
VerneMQ VerneMQ was launched in 2014 by Erlio GmbH. It supports MQTT messages in LevelDB, and uses a clustering architecture based on Plumtree, however it isn’t actively developed and lacks features. NServiceBus NServiceBus is designed with simplicity. With a number of retry strategies, a message which fails processing can automatically be forwarded to an error queue for manual investigation. Kestrel kestrel is a JVM-based distributed message queue inspired by Blaine Cook’s ”Starling”.
In NSQ NSQ , distributed and decentralized topologies are promoted, allowing fault tolerance and high availability as well as reliable delivery by replicating every message across multiple nodes within the cluster. NATS NATS was originally released in 2011 and was written in Go. KubeMQ KubeMQ is a modern and innovative message queue and broker that facilitates communication across cloud platforms, on-premise environments, and edge deployments.
3.2 Proprietary Message Brokers
We found 16 proprietary message brokers, out of which 10 supported priority messages while 6 did not. Each is discussed in more detail in below subsections.
3.2.1 Priority Support
IBM MQ ibmmq supports data exchange between applications, systems, services, and files via messaging queues, serving as a crucial communication layer for message flow management. It offers flexibility in deployment options, whether in virtual machines or containers, including Docker, Kubernetes/Cri-O and Red Hat OpenShift. Moreover, it is ideal for applications demanding high reliability and zero message loss. Amazon Simple Queue Service SQS is operated by Amazon, so it can handle a lot of traffic with providing authentication using the Amazon API key and secret. However, requests are sent to the SQS web service via HTTP, which is susceptible to latency issues.
Microsoft Message Queue MSMQ is a messaging infrastructure created by Microsoft and built into the Windows Operating System. It serves as a queue manager and allows two or more applications to communicate without immediately knowing each other’s responses.
As a Java-based message broker, Oracle GlassFish Server Message Queue  Oracle1 provides message brokering services to popular message queue systems such as AQ, IBM MQ Series, and TIBCO Rendezvous. It provides a consistent, open, JMS-compliant API for these message queuing systems. Additionally, OMB supports both durable and non-durable subscribers, as well as the JMS standard pub/sub, topic-based routing.
TIBCO Rendezvous TIBCORendezvous is a peer-to-peer architecture for high-speed data distribution. TIBCO Enterprise Message Service TIBCOEnterpriseMessageService is a message oriented middleware that supports a wide range of message protocols and technologies, including the Java Message Service (JMS) standard using Java and J2EE, Microsoft .NET, TIBCO FTL, TIBCO Rendezvous and C and COBOL on the Mainframe. Besides supporting up to 10 MB payloads in XML, JSON, CSV, HTML and plain text formats, Anypoint MQ AnypointMQ also has easy connectivity to Mule applications or non-Mule applications.
Azure Service Bus AzureServiceBus from Microsoft is a cloud-based message broker that only supports AMQP and STOMP protocols. As a fundamental part of SAP NW PI SAP architecture, an Integration Broker facilitates communication between different enterprise applications, both SAP-based and non-SAP.
Solace Message Broker Solace , also known as Solace PubSub+, is an advanced event broker that facilitates the efficient exchange of information between applications, IoT devices, and users through several messaging paradigms, including pub/sub, queue, request/reply, and streaming.
3.2.2 No Priority Support
Google Cloud Pub/Sub GoogleCloudPub/Sub is a messaging service offered by Google Cloud. In addition to native integration with other Google Cloud services, including Cloud Functions, Dataflow, and BigQuer, as well as a variety of development tools like Cloud Shell, Cloud Code, and Cloud Build, it supports real-time data processing for ML applications with Google Cloud AI Platform and other ML services. Aside from the Stackdriver Logging and Stackdriver Monitoring tools, it also provides SDKs for Java, Python, Node.js, and Go.
Amazon MQ AmazonMQ developed for ActiveMQ based on Java with support for MQTT, AMQP, STOMP, and WebSocket. Intel MPI Library IntelMPILibrary provides a cloud support for Amazon Web Services, Microsoft Azure, and Google Cloud Platform. Amazon Kinesis Kinesis is a real-time streaming data service with a scalable and durable architecture that can capture and store GBs or TBs of data per second from multiple sources for up to 24 hours. It provides various developer tools and integrations with AWS services, such as SDKs, templates, and integrations with AWS CloudFormation.
Azure Storage Queue AzureStorageQueue provides cloud messaging that enhances communication in the cloud, on desktops, on-premises, and on mobile devices. IronMQ IronMQ runs on public clouds as well as on-premise with providing client libraries in a wide variety of programming languages, including Python, Ruby, Java, PHP, and NET.
3.3 Summary on Message Brokers
Message brokers play a major role in streamlining communication between distributed systems by ensuring messages are properly routed. Among the features which they provide are guaranteed message delivery, queuing of messages, and pub/sub mechanisms. Additionally, they often provide mechanisms for message persistence so that they do not lose a single message in the event of a system failure. Their support for multiple messaging patterns, many different protocols, programming languages, and data styles, meets the needs of various types of communication. Moreover, many brokers come with an array of features. These features are critical elements, each contributing significantly to creating a robust message broker capable of managing communications in complex systems, ensuring efficiency.
Clustering support ActiveMQCluster enables scaling the message service to accommodate more clients or connections, effectively handling large message volumes and numerous clients. Monitoring OpenMQmonitor tools are crucial for tracking a message broker’s performance and health, allowing for proactive management, early problem detection, and reliable operation.
Pub/sub support ibmmq9 enables separation of publishers from subscribers, increasing system flexibility. Parallel processing support Kafka2 allows the message broker to handle multiple messages simultaneously, improving throughput and efficiency.
Pull and push support Kafka2 allow flexible and timely message delivery. Reliable delivery support Oracle5 ensures messages are not lost during transit, typically through acknowledgments, retries, or temporary storage until successful delivery.
Persistence support Oracle6 safeguards messages against loss during broker restarts or storage message failures. Authentication support Oracle7 is vital for permitting only authorized users and systems to post or subscribe to messages, especially in systems dealing with sensitive data.
Scalability RedHat3 refers to a message broker’s ability to handle increasing loads from a large number of concurrently connected clients. Message brokers can operate in distributed environments ibmmq9 , allowing them to work on multiple devices or locations, optimizing workload distribution. Fault tolerance support ibmmq9 enables the message broker to recover from failures and continue operating smoothly.
With these capabilities, they are able to serve as the backbone for a wide variety of event-driven architectures and asynchronous communication patterns by serving as the basis for these architectures.
Despite their inherent benefits, message brokers have several limitations that must be addressed. As well as the difficulties these brokers face in deployment, maintenance, monitoring, integration, and management, they may cause considerable resource consumption. Furthermore, the limited ability of some brokers to handle large amounts of data make them unsuitable for handling real-time applications that are increasingly common in today’s world.
4 Message brokers and GenAI
Figure 3: The overall architecture of a GenAI agent, with possible integration points with message brokers.
Pub/sub communication patterns from the perspective of GenAI hold significant potential, providing a variety of benefits that could enhance content delivery, personalization, and user engagement through a variety of factors. With the use of these AI models, such as GPT-3 and its successors, it is possible to provide customized content in real-time, tailored to individual subscriber preferences by automatically generating data such as human-like text according to individual subscriber preferences wang2023overview .
For example, language-based GenAI systems can process subscriber queries and feedback efficiently using natural language understanding. Moreover, they can power chatbots and virtual assistants that enable users to communicate both interactively and intelligently with each other openai2023gpt4 . Additionally, such systems can sum up content, translate it into different languages, and moderate it, enhancing quality and accessibility garcia2023we .
As a result, systems can be designed to provide valuable information to subscribers proactively in the form of automated reporting, anomaly detection, and predictive analytics 10198233 . By automating content verification and validation, GenAI contributes to scalability, resource optimization, and load balancing wang2023overview .
Moreover, computing architectures are evolving to address the rising processing and performance demands of GenAI applications. The GenAI agent model (Fig. 3) contains four parts including the environment, providing input for the agent’s perception component, the brain component, with decision-making capabilities and LLMs as core computational engines, as well as the actuation component, capable of acting upon the environment through its functional capability to produce actions stuart2010artificial ; xi2023rise . The agent is initialised from the users in the environment, providing predefined goals to the agent.
The perception component perceives the environment through observations. It is equipped with sensing elements which may include physical sensors to gather diverse data from the surrounding environment, software-based tools, or both, along with a message broker designed to facilitate communication between these elements. These physical sensors can capture multi-modal observations, including visual, auditory, and textual data, as well as other modalities that can help the GenAI agent to understand its situation. The software-based tools such as LLMs, interface with abstract data streams. They read, analyze, and transform the gathered data into a comprehensible format for the agent’s brain. These tools includes such as Multimodal-GPT gong2023multimodal , Flamingo alayrac2022flamingo , HuggingGPT shen2023hugginggpt , AudioGPT huang2023audiogpt , GPT-4 openai2023gpt4 , Visual ChatGPT wu2023visual , etc. Using sensors with LLMs enhances their capability to understand and react to changes in the environment, leading to more effective decision-making in dynamic situations.
The agent’s decision-making component, or brain, executes memorizing, thinking, analyzing and decision-making tasks, supporting long and short-term memory, knowledge, and planning xi2023rise . Short-term memory records recent tasks and actions, while long-term memory acts as an external database, enhancing the agent’s ability to recall past conversations and pertinent details. Utilizing subgoal decomposition wang2023survey and a chain-of-thought approach wei2022chain , the agent breaks down large tasks into multiple manageable subgoals that are processed by a group of LLMs models. Through self-critics and reflection yao2022react ; shinn2023reflexion , the agent can learn from its errors, enhancing its capabilities iteratively.
The agent uses actual physical actuators, LLM- based tools, or both to execute tasks in the actuation component. These elements allow agents to interact with and respond to their environments. The LLM-based tools include text generation through text-based tools such as ChatGPT chatgpt2023 . Moreover, agents’ workspaces have been expanded with embodied actions to support their integration and interaction with the physical world. LM-Nav shah2023lm analyzes input commands and the environment, aiming to identify the optimal walk based on a topological graph that is constructed internally. EmbodiedGPT mu2023embodiedgpt enables robots to comprehend and perform motion sequences in physical settings through multimodal visual understanding. Using these non-textual output tools extend the functionality of language models and agent scenarios.
Crucially, GenAI agents are also able to generate novel tools. With frameworks like CREATOR qian2023creator , agents can generate executable programs or merge existing tools into more robust ones. Furthermore, with frameworks such as SELF-DEBUGGING chen2023teaching , agents can iteratively improve the quality of the generated tools, autonomously learning from past experience, self-correcting and adapting, enhancing their tool-generation capabilities.
Following milestone studies in edge intelligence park2019wireless ; loven2019edgeai ; deng2020edge , the interaction between brokers and GenAI can be separated into two different categories: the benefits GenAI can bring for message brokers, and the benefits message brokers can bring for GenAI. Fig. 4 further illustrates the expected benefits of integrating a message broker with GenAI.
Figure 4: Benefits provided by integrating a message broker with a GenAI agent.
4.1 GenAI for message brokers
GenAI can enhance message broker intelligence and efficiency, specifically through predicting routing decisions (by avoiding busy routes) and filtering of messages tailored to specific criteria dhoni2023exploring . Moreover, using GenAI’s predictive capabilities dhoni2023exploring , the message broker can proactively distribute workloads or services across various nodes with minimal access times.
GenAI’s ability to identify error-inducing patterns dhoni2023exploring enables the message broker to implement automatic corrective measures, increasing system reliability and reducing downtime.
Additionally, GenAI can determine optimal times to scale resources so that resources are utilized optimally dhoni2023exploring . Furthermore, GenAI’s ability to learn from ongoing interactions and explainability results in a continuous feedback loop openai2023 that refines system performance.
4.2 GenAI on message brokers
By contrast, asynchronous communication capability ibmmq9 allows message brokers to significantly improve GenAI task processing efficiency by facilitating interactions that do not require real-time communication. This setup also streamlines real-time data stream processing.
A key component of message brokers is their ability to prevent message loss Oracle4 , which is essential for GenAI to prevent misuse of computational resources or loss of critical data. Moreover, message brokers often feature tools that monitor and manage the flow of messages, a key element of maintaining performance and operational integrity of GenAI.
Furthermore, message brokers facilitate service decoupling ibmmq9 , enabling GenAI to manage growing workloads more effectively. The broker’s ability to distribute tasks to different nodes allows GenAI to achieve a balanced load distribution ibmmq9 , with different components running on different nodes. This is particularly important for GenAI deployments in the computing continuum, providing them with access to local environments with limited computational capacity. In such cases, a message broker can act as an intermediary, bridging sensors with the perception module and conveying actions to the actuators and responses to the user.
Moreover, within in the brain component, a message broker can connect the different components, which often require heavy computational capacity. Fostering collaboration between the different LLM models forming up the decision-making component, message brokers can improve agent performance on complex tasks and allow the offloading of tasks in a balanced way among the available resources, thus enabling for efficient use of computing, communication and storage resources.
However, integrating GenAI into the pub/sub form requires careful planning and customization. GenAI may lead to biased and misleading outcomes. Thus, it is crucial to address privacy and ethical issues while aligning with specific system goals 10198233 . In particular, message brokers need advanced methods to handle the massive amounts of data potentially consumed and generated by GenAI applications.
As an example, while 5G or 6G networks can increase connectivity, model compression methods and model training acceleration may reduce GenAI computational requirements and delays. Furthermore, the efficient resource management process enables workload distribution based on the capacity of each node, thus improving the performance of the distributed system. Moreover, by turning GenAI input and output data into priority-based smart data, timely processing can be further improved. Furthermore, choosing the right models and updating them as data changes improves results. Finally, a continuous monitoring system can detect anomalies and data loss immediately, allowing for corrections and adaption in real time.
In the following sections, we will discuss the advanced methods designed to enhance message brokers’ functionality within the GenAI context, overcoming deployment challenges. By tackling these issues, we aim for seamless integration and optimal performance of GenAI applications.
4.3 Semantic communication
Effective management and real-time analysis of vast data volumes are crucial for developing GenAI applications. This process must be flexible, accommodating various data types with utmost precision. To alleviate strain on communication networks, it’s essential to embed efficient (e.g., semantic) communication mechanism xia2023generative within message brokers. Such integration can enable intelligent, automated feature selection aligned with subscriber needs, streamlining the transmission of data-rich content donta2023governance . Consequently, this approach facilitates smart data transfer, optimizing resource allocation and minimizing latency.
Message brokers equipped with dynamic prioritization capabilities can intelligently identify and route high-priority messages by incorporating semantic communication technology. Prioritization allows message brokers to handle messages based on content along with criteria within message header. This feature not only reduces network congestion but also preserves network resources. Additionally, integrating semantic communication with a broker requires high scalability, processing power, and memory to handle big data, while maintaining robust security features to process sensitive data.
In delay-sensitive applications like healthcare, this integration is vital to assigning priorities based on its deep understanding of data patterns and sensitivities and its subscribers’ specific needs. For example, in the healthcare scenario, GenAI can analyze large amounts of medical data to identify urgent cases, alerting healthcare professionals of critical patient needs or alarming health trends oniani2023military . Similarly, GenAI can prioritize alerts based on the severity and location of incidents at fire stations, ensuring a prompt response.
As examples, integration of semantic communication with RabbitMQ improves message routing capabilities through embedding semantic annotations into messages. As a result of these annotations, the broker can understand the meaning and context of each message, enhances message routing and efficient delivery.
Furthermore, Apache RocketMQ, with its support for message broadcasting and filtering ApacheRocketMQdocs , has the potential to offer advanced message routing and filtering capabilities by leveraging semantic communication. These capabilities involve interpreting messages’ semantic content, allowing for more precise and context-sensitive handling.
Finally, Anypoint MQ’s intelligent routing AnypointMQ1 combined with semantic filtering allows messages to be identified according to their inherent importance and meaning. It also ensures faster processing, routing of urgent and high priority content, and timely responses.
4.4 Dynamic data and model management
GenAI-based applications require a data and model management system that not only simplifies the construction of AI models but also optimizes efficiency and effectiveness. Such a system should minimize the need for human intervention in selecting ML models, enhancing real-time responsiveness with high model’s accuracy, and improving real-time inference capabilities when integrated with message brokers. LLMs can expedite the model selection process and boost the deployment efficiency and precision of AI solutions zhang2023mlcopilot ; shen2023large .
Moreover, integrating GenAI models with message brokers involves managing and directing the related data flows. Since processing GenAI models is computationally intensive, with large amounts of real-time data, it needs to be scalable and distributed to cope with the varying workloads. Such integration also must also consider the handling of streaming data.
As an example, due to its distributed streaming capabilities, Apache Pulsar Pulsar can be integrated with GenAI models that often require handling large datasets for training and inference. Brokers also need to handle a variety of data formats for integration with GenAI models. Red Hat AMQ, supporting multiple message patterns RedHatAMQ , could ensure compatibility with diverse GenAI data types.
Because GenAI models such as LLMs often process sensitive or personal data, robust security measures are critical. For example, Google Cloud Pub/Sub GoogleCloudPub/Sub incorporates multiple integrated security measures to protect confidential data being transmitted. These measures include authentication and encryption using keys managed by Google, along with the latest Data Encryption Key (DEK) technology, among others GoogleCloudPub/Sub6 .
Moreover, Amazon Kinesis Kinesis offers reliable data transmission methods, utilizing synchronous data replication, checkpointing mechanisms, and error handling strategies Kinesis2 ; Kinesis3 .
Further, Apache Kafka’s Kafka robust architecture, equipped with a stream processing feature, a cluster of brokers, and topic partition features, enables it to efficiently handle parallel processing tasks and feed the data to GenAI models.
4.5 Training acceleration
The training of GenAI models requires a significant amount of computation and time. By incorporating training acceleration methods li2023colossal into a message broker system, training time can be reduced, computational resources saved, and models deployed more rapidly.
The most common technique is sequence parallelism (SP), in which the input sequence is divided into smaller sub-sequences and processed in parallel li2021sequence . Another method involves selective activation re-computation, in which only the necessary parts of the model are recalculated during training, rather than the entire model korthikanti2023reducing . A fine-tuning technique involves adjusting the parameters of an existing model rather than training a new one from scratch. As a result, training data and computations can be reduced radford2018improving .
Furthermore, tensor parallelism works by splitting the model across multiple GPUs and processing different parts of the model in parallel li2023colossal . The mixed-precision training technique is one of the most effective ways to speed up training on modern GPUs micikevicius2018mixed . In this technique, the amount of memory required during training is reduced by using lower-precision data types.
As of our knowledge, no broker currently supports model training acceleration. However, a message broker offers a variety of useful functionalities, including the scheduling and distribution of data compression tasks, as well as monitoring the performance of compressed models. These methods are discussed in more detail in Sections 4.7 and 4.8.
4.6 Dynamic Model Compression
Integrating GenAI models on resource-constrained nodes in the computing continuum often requires model compression, especially with high-dimensional models, strenuous computational tasks, and low latency requirements bariah2023large . Among the compression strategies, pruning is prominent, involving the removal of superfluous elements from a model to decrease its size and complexity jiang2022model . Another technique is Knowledge Distillation, where a smaller ’student’ model learns to replicate the functionality of a larger ’teacher’ model zhu2023survey ; polino2018model .
Furthermore, various quantization methods are utilized to decrease model parameter precision, thereby cutting memory usage and computational needs substantially polino2018model . Another technique is low-rank factorization, which simplifies weight matrices with lower-rank approximations to reduce model size and computational demands hsu2022language .
Currently, no broker supports model compression as such. However, message brokers can provide essential capabilities such as scheduling and distributing compression tasks, or monitoring the performance of compressed models. These methods are discussed in more detail in Sections 4.7 and 4.8.
4.7 Dynamic Orchestration
Integration with GenAI requires effective resource management. The broker must allocate and manage resources such as CPU, memory, and storage to handle computational requirements. Furthermore, supporting feedback mechanisms is essential to allow continuous learning and improvement of the system based on the output and performance of the GenAI agent. However, the process of efficient resource usage and allocation requires parallel processing capability.
For instance, Celery can handle, maintain, and schedule distributed tasks across various nodes, thereby enhancing the efficiency of GenAI agents Celerydev ; Celeryintro . Apache ActiveMQ provides efficient resource allocation ApacheActiveMQ . Thus, optimizing Apache ActiveMQ is critical to providing intelligent resource allocation to ensure balanced workloads across different nodes.
Amazon SQS provides a reliable queue service for handling messages that facilitates microservices and distributed systems decoupling SQS1 . Further, the robust architecture of Apache Kafka, which uses clusters of brokers to handle data distribution and partition topics for scalability, makes it excellent for supporting distributed event streaming Kafka1 . Additionally, Azure Service Bus provides advanced scheduling features and message orchestration in distributed environments AzureServiceBus5 ; AzureServiceBus4 .
4.8 Dynamic System Monitoring
MLOps, merging DevOps principles with machine learning (ML), is central to the advancement of ML and AI, streamlining the lifecycle of models from development to deployment. A critical feature within this field is the monitoring of deployed models, crucial for the uninterrupted and reliable operation of message broker systems. This practice enables real-time insights into model metrics, resource usage, and system irregularities, creating a proactive environment for identifying and addressing issues promptly. Furthermore, MLOps facilitates the setting up of automated alerts and triggers, enhancing responsiveness to anomalies and minimizing downtime loven2023can ; kreuzberger2023machine .
A Continuous Diagnostics and Mitigation (CDM) program plays a vital role in network security by analyzing network behavior and thwarting unauthorized access, thereby enabling prompt responses and maintaining network integrity. Beyond autonomous configuration management and monitoring device availability, CDM programs conduct continuous health assessments of devices and evaluate their environmental footprint. This continuous surveillance helps identify potential threats, bolstering processes to enhance security measures. Furthermore, CDM ensures the protection of sensitive information against unauthorized access or breaches Praveen2023exploring .
A message broker can be embedded with monitoring tools like MLOps and CDM, but it is a complex process and resource hungry. It requires distributed architectures and parallel processing to handle high-intensive tasks effectively and timely. There is a chance of starvation between regular data transmissions and monitoring. However, this integration demands the message broker’s ability to allocate and manage computational resources intelligently. This requires striking an optimal balance between monitoring and regular tasks to gain performance optimization, lower cost, and energy efficiency.
As KubeMQ supports multistage pipelines KubeMQ2 , it could be essential for managing data pipelines within MLOps frameworks. Moreover, HiveMQ and Amazon Kinesis can contribute to MLOps integration by facilitating real-time monitoring and alert systems integration, since they support real-time device monitoring HiveMQ ; Kinesis .
Finally, tracking GenAI model performance can be effectively achieved through integration with IBM MQ, which has robust monitoring and tracing capabilities ibmmq2 . Solace PubSub could potentially assist in identifying and resolving issues related to message routing, delivery, and processing Solace1 , thus facilitating control and modeling activities in MLOps environments.
4.9 Impact of 5G and 6G on message brokers
The advent of 5G technology has significantly influenced message brokers by enabling faster data rates, reduced latency, and enhanced connectivity and reliability. 5G efficiently manages vast message volumes, crucial for real-time applications, and supports numerous devices simultaneously without triggering network congestion 8767268 .
The anticipated 6G evolution promises further enhancements, such as higher throughput, even lower latency, and AI integration at the network level loven2019edgeai ; peltonen20206g ; peltonen2022many . These advancements will empower message brokers to dynamically optimize message routing, adapt to fluctuating network conditions, assure consistent reliability, and process extraordinary data volumes. This capacity is essential for more sophisticated applications necessitating stringent data security. To that end, 6G is projected to fortify security protocols, a critical aspect of protecting future message broker systems porambage2019sec .
5 Conclusion
In this paper, we summarized recent message brokers by highlighting their features and limitations. We categorized these message brokers using criteria such as their open-source availability, integrated monitoring tools, message prioritization mechanisms, parallel processing capabilities, reliability, clustering and distribution capabilities, authentication processes, data persistence strategies, fault tolerance, and scalability.
Further, we explored the limitations of the message brokers, and reflected on requirements for a robust message broker framework that would meet GenAI specifications. To that end, we discussed possible approaches for enhancing the effectiveness of message brokers to support the rapid development of GenAI applications. These approaches include, for example, model training acceleration, model compression, resources orchestration, large-scale data processing, and monitoring systems performance. Furthermore, we discussed the impact of 5G and 6G technologies on message brokers.
By examining the features, challenges, and potential approaches for future research on message brokers, our study will provide the basis for developing a flexible communication system that can distribute data efficiently, provide high-quality service, manage resources intelligently, and meet the needs of ever-growing GenAI applications.
Table 1: A summary of open source and priority-supporting message brokers.
Message
Brokers and Queues
Clustering Support
Monitoring Support
Pub/Sub Support
Parallel Processing
Pull & Push Support
Reliable Delivery
Persistent
Authentication
Scalable
Distributed
Fault Tolerance
Shortcomings
Features
Apache ActiveMQ ApacheActiveMQ Both
Message delivery guarantees are limited ActiveMQDocument .
Memory per queue is limited, the default number
of messages is 400 ActiveMQPolicies .
Installation is complex ApacheActiveMQ .
Scaling is challenging ActiveMQPolicies .
Multi-protocols & multi-languages support ApacheActiveMQ .
Efficient management and resource allocation ApacheActiveMQ .
Supports flow control and message expiration ApacheActiveMQ .
Provides message groups as well as virtual and combined queues ApacheActiveMQ .
Works on small and medium-scale applications ActiveMQPolicies .
Fuse Message Broker Fuse Both
Limited monitoring toolsFuse1 .
Written in Java Fuse1 .
Supports JMS 1.1 & J2EE 1.4 integration-related components Fuse1 .
Supports loosely couple applications Fuse1 .
Supports multi-languages including C/C++, Java, .NET, Ruby, Perl,
PHP, Pike,& Python Fuse1 .
Supports message compression Fuse1 .
Apache Qpid ApacheQpid Both
Compatibility issues between versions ApacheQpidReleases .
Message size is limited to 100MB for AMQP
protocols 0-8, 0-9, or 0-91 ApacheQpidsize
Implements AMQP Protocol ApacheQpid .
Easy to use ApacheQpid2 .
Detects failures and assigns messages to different brokers ApacheQpidConn .
Low latency ApacheQpid3 .
Supports multiple authentication schemes ApacheQpid1 .
Active connections can be limited to protect client processes
from malicious activity ApacheQpidbroker .
RabbitMQ RabbitMQ Push
Written in Erlang, which is unfamiliar to many
developers RabbitMQUser .
Queues with large numbers of messages are memory-
intensive and strain brokers RabbitMQcmemory .
Redundant message broker communication donta2022survey .
Clustering has few features and is complicated RabbitMQcheck .
Message size is limited to 512MB donta2022survey ; RabbitMQcons .
Runs on all major operating systems RabbitMQplatforms .
Has good documentation RabbitMQdocs .
Works with C, C++, .NET, and Python RabbitMQtools .
Supports asynchronous cluster-to-cluster message routing RabbitMQ1 .
Supports multiple messaging protocols RabbitMQprotocol .
Offers several built-in exchange types RabbitMQfeature .
Supports flow control for balancing workloads and avoiding rapid
messages flooding RabbitMQflow .
HornetQ HornetQ Both
Data loss may occur donta2022survey .
Delay may occur with large messages (up to 100KB)
due to split message into multiple packages HornetQuser .
Supports AMQP and STOMP protocols HornetQprotocol .
Provides better performance and stability when combined with
ActiveMQ donta2022survey .
Red Hat AMQ RedHatAMQ Both
Queue access is limited by special characters RedHatAMQ1 .
Non-persistent messages are lost when brokers
stop RedHatAMQ2 .
Enables real-time integration RedHatAMQ .
Supports multi-message patterns for real-time messaging RedHatAMQ .
Supports multi-languages, including Java, C, C++, Python, Ruby,
& .Net RedHatAMQ .
Supports mission-critical applications RedHatAMQ .
Celery Celery Push
Compatibility and integration with other brokers
can be complicated Celeryintro .
Overall complexity Celeryintro .
Monitoring and management are challenging Celerymonitor .
Number of connections is limited by 10 connections Celerylimit .
Enables operations to manage and maintain distributed task queues
,such as starting, stopping, and restarting worker processes Celerydev .
Functions as a task queue Celeryintro .
Focuses on real-time processing Celerydev .
Supports task scheduling Celeryintro .
Supports multi-message brokers Celeryintro .
Integrates with multi-web frameworks Celeryintro .
Supports automatic retry in the event of connection loss or
failure Celerytask .
JBoss Messaging JBossMessaging Push
Delay may occur with large messages (up to 100KB) due
to split message into multiple packagesJBossCons .
Supports AMQP, MQTT, STOMP message protocols JBosspro .
Supports transactions JBosstra .
Provides management processes related to deployments, configuration,
and access control JBosscon .
Easy integration with other JBoss and Java $E$ components JBoss31 .
OpenMQ OpenMQ Both
Setup is complex OpenMQoverview .
High Latency donta2022survey .
Loosely-coupled architecture OpenMQoverview .
Beanstalk Beanstalk Pull
Lacks authentication Beanstalk .
Message size is limited to 64 KB Beanstalkprotocol2 .
Unprocessed messages are automatically returned to the queue Beanstalkprotocol2 .
Supports Ruby, Rails, Java, JavaScript, Haskell, and PHP Beanstalkprotocol .
Gearman Gearman Both
Does not have authentication and SSL support Gearman .
Manual configuration Gearmandoc .
Monitoring tools are limited Gearmanprotocol .
Used by LiveJournal, Yahoo!, and Digg GearmanIntro .
Multi-languages support Gearmandoc .
No single point of failure Gearman .
No limits on message size Gearman .
Supports load balancing Gearman .
Enduro/X Enduro/X Both
Message size is limited to max 10 MB Mavimax2022Enduro/X1 .
Buffer size is limited to max 64KB Enduro/X1 ; Mavimax2022Enduro/X1 .
Cluster nodes number is limited to max 32 nodes Mavimax2022Enduro/X .
Resource managers numbers with single transaction
are limited to max 32 Mavimax2022Enduro/X .
Versions compatibility depends on the date of
release Enduro/X2 .
Limitations on availability of the operations that can be
executed within the callback Enduro/X1 .
Distributed transaction processing Mavimax2022Enduro/X .
Works on multi-platforms Enduro/X .
WSO2 WSO2 Both
Heap memory size allocation is limited to max
4GB WSO2memory .
Supports widely used protocols such as HTTP/S, JMS, VFS, UDP,
TCP, MQTT, MSMQ, and MailTo WSO2transport .
Supports message filtering WSO2filter .
Integrates easily with other WSO2 products and third-party
systems WSO2 .
HiveMQ HiveMQ Push
Number of characters broker accepts in an Client ID is
limited between 1 and 65535 HiveMQ2 .
Number of characters broker accepts in a topic string is
limited between 1 and 65535 HiveMQ2 .
Resource intensive for maintenance HiveMQ3 .
Is a client-based MQTT broker for M2M communication HiveMQ .
Suitable for mission-critical applications HiveMQ1 .
Supports real-time monitoring of device data & integration with
existing systems HiveMQ .
Redis Redis Push
Uses a memory dump which leads to slow
performance Redis1 .
Has only basic security options Redis2 .
Supports multiple data types Redis .
In-memory data storage Redis .
EMQX EMQX Both
Setup, configuration and management are complex EMQX4 .
Supports MQTT bridging EMQX2 .
Supports data integration EMQX1 .
Apache Pulsar Pulsar Push
Complex configuration and deployment Pulsar2 .
Complex architecture, based on four components (Pulsar
servers, Apache BookKeeper, Apache ZooKeeper, and the
RocksDB database) that need to be configured and
managed Pulsar1 .
Supports event streaming Pulsar .
An index-based storage system Pulsar1 .
Low latency Pulsar .
Supports messaging, streaming, and queuing Pulsar .
Table 2: A summary of open source and non-priority-supporting message brokers.
Message
Brokers and Queues
Clustering Support
Monitoring Support
Pub/Sub Support
Parallel Processing
Pull & Push Support
Reliable Delivery
Persistent
Authentication
Scalable
Distributed
Fault Tolerance
Shortcomings
Features
Apache Kafka Kafka Pull
Resource intensive Kafka1 .
Provides a data backlog Kafka1 .
Complex Kafka1 .
Supports topic (log) compaction & distributed event streaming Kafka1 .
Supports data integration Kafka .
Language support Kafka .
Supports multiple data formats Kafka1 .
Supports permanent storage & the management of data flow and
consumer groups Kafka1 .
Supports replication and partitioning of data Kafka1 .
Supports deployment in different environments Kafka1 .
Apache RocketMQ ApacheRocketMQ Both
Message size is limited to max 4MB ApacheRocketMQ2 .
Message sending retries is limited to max 3 times ApacheRocketMQ2 .
Supports message broadcasting, tracking, filtering, and retrying ApacheRocketMQdocs .
Low latency ApacheRocketMQ1 .
Maintains the order of messages ApacheRocketMQdocs .
Supports multi-protocols ApacheRocketMQoverview .
Supports multiple programming languages ApacheRocketMQoverview .
Eclipse Mosquitto EclipseMosquitto Both
Limited security EclipseMosquitto1 .
No built-in clustering EclipseMosquitto6 .
Unsuitable for large-scale deployments EclipseMosquitto6 .
Deployment is challenging in a cloud environment EclipseMosquitto6 .
Message size is limited to max 256MB EclipseMosquitto .
Low resource usage EclipseMosquitto3 .
QoS support EclipseMosquitto4 .
Topic-based message filtering EclipseMosquitto5 .
Supports logging and debugging EclipseMosquitto4 .
Supports functioning as a bridge EclipseMosquitto4 .
Supports dynamic restart configuration EclipseMosquitto4 .
Suitable for low-power machines EclipseMosquitto3 .
ZeroMQ zeromq Both
High load of local control modules donta2022survey .
Fails to manage relationships between all network
components donta2022survey .
Limited security zeromq3 .
Scaling is challenging zeromq2 .
Delivery is not guaranteed zeromq2 .
Brokerless messaging platform zeromq1 .
Multi-languages and platforms support zeromq .
Carries messages across IPC, TCP, TPIC, and multicast zeromq1 .
Low latency zeromq .
Apache NiFi ApacheNiFi Both
Data extraction is difficult when a node is
separated from a cluster ApacheNiFidoc .
Under certain conditions, data is automatically
deleted ApacheNiFidoc .
Complex configuration ApacheNiFiConfigure .
Provides a data flow framework ApacheNiFidoc .
Provides data compression using a user-specified algorithm to reduce
data size ApacheNiFidoc .
Prevents data loss by controlling data flow and stopping the
production of more data than a queue can handle ApacheNiFidoc .
Supports buffering of all queued data ApacheNiFioverview .
Integrates and processes multiple data sources ApacheNiFioverview .
Ably Realtime AblyRealtime Both
Caps the number of channels per connection AblyRealtime2 .
Peak connections number limited between 200 and
240 AblyRealtime2 .
Message size limited to 16KB AblyRealtime2 .
Number of queues limited to 5 AblyRealtime2 .
Queue length limited to 10,000 AblyRealtime2 .
Addresses challenging real-time requirements AblyRealtime .
Supports streaming data AblyRealtime1 .
Supports multiple protocols AblyRealtime .
Apache SamZa SamZa Pull
Only supports JVM languages SamZa2 .
Configuration is complex SamZa .
Resources intensive with large data volumes kleppmann2019apache .
Stream processing framework SamZa .
Supports message storage, routing, and processing management SamZa .
Supports at-least once data processing SamZa1 .
Real-time data processing with low latency SamZa1 .
Easy to integrate SamZa1 .
VerneMQ VerneMQ Both
Lack of security donta2022survey .
Clustering architecture is unproofed EMQX5 .
Limited enterprise features EMQX5 .
Not under active development EMQX5 .
Limited support for MQTT integration EMQX5 .
Lacks management and monitoring features EMQX5 .
No cloud-based service VerneMQ .
Master-less clustered messaging protocol VerneMQ1 .
Supports flow control donta2022survey .
Low latency VerneMQ .
NServiceBus NServiceBus -
Scalability is limited due to use centralized
resource NServiceBus3 .
Monitoring tools are limited NServiceBus4 .
Debugging is complex with huge stream of
messages NServiceBus5 .
Ensures message processing NServiceBus .
Supports transactions and recovery is built-in NServiceBus2 .
Messages can be retried at regular intervals NServiceBus1 .
Kestrel kestrel Pull
Low support of security kestrel .
Low clustering capabilities kestrel1 .
Memory size is limited to max 128MB kestrel .
Number of items in the queue is limited to 500 kestrel1 .
Data size of each item in the queue is limited to max
32bytes kestrel2 .
Written in Scala kestrel .
Each server handles ordered MQs, with no cross
communication, resulting in a cluster of k-ordered queues kestrel .
NSQ NSQ Push
Data loss with server crash NSQ2 .
Messages are unordered NSQ2 .
Limited persistence NSQ2 .
No message recovery NSQ2 .
Lacks replication NSQ2 .
Messages are delivered at least once, which may
duplicate messages NSQ2 .
Load-balanced message delivery NSQ .
Efficient handling of high-volume& real-time data streams NSQ ; NSQ1 .
NATS NATS Both Message size is limited to max 64MB NATS2 .
Suitable for real-time communication NATS1 .
Easy to use NATS1 .
Minimal resource consumption NATS1 .
Offers persistence with ”at-least-once” and ”exactly-once” NATS2 .
KubeMQ KubeMQ Both
Unsuitable to all use cases due to it’s designed for
dynamic microservice environments ayaz2022data .
Builds a hybrid infrastructure across clouds, on-prem, and at the
edge to allow microservices from multi-environments to
communicate KubeMQ .
Support for pub/sub, microservices, multistage pipeline,
and tasks queue use cases KubeMQ2 .
Runs in Kubernetes and connects natively to the K8S
cloud-native ecosystem KubeMQ .
Simple deployment in Kubernetes KubeMQ .
Easy to use KubeMQ .
Low latency KubeMQ3 .
Table 3: Summary of proprietary and priority-supporting message brokers.
Message
Brokers and Queues
Clustering Support
Monitoring Support
Pub/Sub Support
Parallel Processing
Pull & Push Support
Reliable Delivery
Persistent
Authentication
Scalable
Distributed
Fault Tolerance
Shortcomings
Features
IBM MQ ibmmq Push
High costs ibmmq1 .
Problems with message prioritization due to unordered way
of allocating messages IronMQibm .
Does not always integrate with the newest forms of
messaging IronMQibm .
Messages are unordered IronMQibm .
QoS support ibmmq8 .
Provides robust monitoring & tracing of all
messages ibmmq2 .
Controls undelivered messages ibmmq4 .
Multi-APIs support ibmmq5 .
Allows applications to be decoupled ibmmq3 .
Easy to deploy on various platforms ibmmq3 .
Amazon SQS SQS Pull
High scale-up cost SQS2 .
Message size is limited between 1KB and 256 KB SQS3 .
Message ordering is not guaranteed SQS4 .
Message retention before deletion is limited between 1 minute
and 14 days SQS4 .
Cloud-based web service SQS1 .
Supports decoupling microservices, distributed
systems, and serverless applications SQS1 .
Transmits, stores, and receives messages across
software components using SQS at any volume SQS .
Messages are delivered at least once SQS4 .
Microsoft MQ (MSMQ) MSMQ Push
May experience resource failure MSMQ6 .
Limitation on message size MSMQ6 .
Drops all MSMQ messages if the appropriate server is not
deployed MSMQ7 .
Open queue failure error prevents data transfer MSMQ7 .
Multi-protocols support MSMQ1 .
Tracks and deletes expired messages MSMQ5 .
Manages distributed brokers MSMQ4 .
Supports remote access MSMQ4 .
Effective routing MSMQ2 .
Provides guaranteed message delivery MSMQ2 .
Provides a store and forward mechanism MSMQ4 .
Supports transactions MSMQ3 .
Oracle GlassFish Server
Message Queue Oracle1
Both
Resource intensive as the number of messages increases Oracle3 .
High latency as connections number to the broker
increases Oracle3 .
Size of message is limited to max 70MB Oracle1 .
Supports transactions Oracle1 .
Supports JMS 1.1, STOMP, and HTTP
protocolsOracle1 .
Well-known standards-based messaging
support Oracle1 .
TIBCO
Enterprise
Message Service TIBCOEnterpriseMessageService
Both
Excludes fault tolerance of the server TIBCO2 .
Recourse intensive as message size increased up to 512MB TIBCO3 .
Supports load balancing TIBCO1 .
Manages the real-time flow of information TIBCOEnterpriseMessageService .
Supports multiple message protocols and
technologies TIBCO1 .
Easy integration with TIBCO eFTL™ software
expands broker to web and mobile
applications TIBCOEnterpriseMessageService .
Loosely coupled design TIBCOEnterpriseMessageService .
Supports integration for heterogeneous
platforms TIBCOEnterpriseMessageService .
TIBCO Rendezvous TIBCORendezvous Both
Expensive TIBCORendezvous3 .
Queue size limited to max 500 TIBCORendezvous2 .
Supports C, C++, Java , & .NET programming
language TIBCORendezvous1 .
Easy to use & setup TIBCORendezvous5 ; TIBCORendezvous4 .
Has a distributed architecture to eliminate
failure TIBCORendezvous1 .
Anypoint MQ AnypointMQ Pull
Expensive AnypointMQ3 .
Payload size is limited to max 10 MB AnypointMQ2 .
Converts the payload format, leading to an increase in
payload size. AnypointMQ5 .
Supports data integration AnypointMQ4 .
Stores messages in a queue AnypointMQ .
Provides intelligent message routing AnypointMQ1 .
Azure Service Bus AzureServiceBus Both
Is a cloud-only service AzureServiceBus .
Message size is limited between 256 KB and 100 MB AzureServiceBus3 ; Azure .
Number of queues is to max 10,000 Azure .
Number of subscriptions per topic is limited to max 2,000 AzureServiceBus3 .
Provides duplicate detection, duplicate
messages will not be stored in the queue AzureServiceBus2 ; Azure .
Guarantees ordering Azure .
Offers scheduling AzureServiceBus5 ; AzureServiceBus4 .
Integrates well with other Azure products AzureServiceBus4 .
Supports multi-protocols AzureServiceBus1 .
Provides delivery guarantee (at-least-once,
at-most-once) Azure .
SAP NW PI SAP Both
Message size is limited to max 350 MB SAP3 .
Performance is directly affected by the message size SAP3 .
Supports various integration patterns SAP1 .
Supports message transformation SAP2 .
Solace PubSub Solace Both
Message size is limited to 64 MB Solace5 .
Broker connections is limited to max 1000, some functions
are not available with the default 100 connections Solace4 .
Guaranteed messaging is not supported over connections Solace3 .
Provides dynamic message routing Solace1 .
High availability & high-performance Solace1 .
Provides distributed tracing Solace2 .
Event-driven architecture Solace1 .
Supports multi-protocols Solace1 .
Table 4: Summary of proprietary and non-priority-supporting message brokers.
Message
Brokers and Queues
Clustering Support
Monitoring Support
Pub/Sub Support
Parallel Processing
Pull & Push Support
Reliable Delivery
Persistent
Authentication
Scalable
Distributed
Fault Tolerance
Shortcomings
Features
Google Cloud
Pub/Sub GoogleCloudPub/Sub
Both
Unsuitable for large-scale deployments
due to limitations on resources GoogleCloudPub/Sub5 .
Message size is limited to max 10MB GoogleCloudPub/Sub5 .
Provides real-time stream analytic GoogleCloudPub/Sub .
Handles the underlying infrastructure,
including provisioning servers, monitoring,
scaling, backups, and security updates GoogleCloudPub/Sub1 .
Provides service maintenance feature that suitable
for Google’s most fundamental products to serve
all customers effectively GoogleCloudPub/Sub2 .
Provides system maintenance feature to detect
any issues with releases by continuously-running
tests it is before used by customers and by
monitoring GoogleCloudPub/Sub2 .
Integrates with other Google Cloud services GoogleCloudPub/Sub .
Supports automatic retries and message
ordering GoogleCloudPub/Sub3 ; GoogleCloudPub/Sub4 .
Azure Storage Queue AzureStorageQueue Pull
Orders messages randomly Azure .
Message size is limited to max 64 KB AzureStorageQueue .
Maximum number of queues is unlimited Azure .
Activity monitoring support Azuremonitor .
Supports storing large numbers of messages AzureStorageQueue .
Messages are delivered at-least-once Azure .
Does not provide duplicate detection Azure .
Amazon MQ AmazonMQ Both
Number of broker connections
is limited to 1,000, or 100 for micro
brokers AmazonMQ2 .
Supports Standard Java Message Service (JMS)
features AmazonMQ2 .
Performs maintenance to the hardware,
operating system,& the engine software
a message broker AmazonMQ3 .
Integrates with other AWS services and
applications AmazonMQ1 .
Supports distributed transactions AmazonMQ2 .
Multi-protocols support AmazonMQ2 .
Intel MPI Library IntelMPILibrary Both
Expensive IntelMPILibrary2 .
Compatibility issues with some
systems IntelMPILibrary ; IntelMPILibrary1 .
Other processor architectures are not
supported IntelMPILibrary .
Uses OpenFabrics Interface (OFI) to handle all
communications IntelMPILibrary .
Establishes the connection only when needed,
which reduces the memory footprint IntelMPILibrary .
Chooses the fastest transport available IntelMPILibrary .
Supports multi-cloud platforms IntelMPILibrary .
Supports multi-cluster interconnects IntelMPILibrary .
Amazon Kinesis Kinesis Push
Permission issues Kinesis5 .
Costly as data volume increases Kinesis4 .
Data payload size is limited to max
1MB), data read rate to to 1MB/s,
and number of consumers for each
data stream to max 20 Kinesis6 .
Operations are rate-limited Kinesis6 .
Limitation on number of data
streams Kinesis7 .
Each record is added to a buffer with a
deadline Kinesis8 .
Provides buffering and processing of real-time
data streaming Kinesis .
Serverless streaming data service Kinesis .
Provides reliable data processing and delivery with
checkpointing and error-handling Kinesis ; Kinesis3 .
Offers monitoring and management Kinesis .
Offers various developer tools Kinesis1 .
IronMQ IronMQ Both
Expensive IronMQ2 .
Limited control IronMQ1 .
Meets the needs of both small businesses and large
enterprises IronMQ1 .
Supports multiple programming languages IronMQ1 .
Uses REST API IronMQ .
Easy to install IronMQ1 .
Handles load buffering, synchronicity, and database
offloading issues IronMQ1 .
No limitation on the number of queues IronMQ .
Acknowledgement
This research is supported by the Research Council of Finland through the 6G Flagship program (Grant 318927), and by Business Finland through the Neural pub/sub research project (diary number 8754/31/2022).